Loss incurred in Day 3 of trading = 5,000 bushels ✕ 3 ✕ $0.01

= $150

Cumulative Profits remaining after loss incurred

= Profits from day 2 - Loss incurred in Day 3

= $450 - $150

= $300

Amount in margin account balance at the end of Day 3 trading

= Value remaining after loss incurred + Initial Margin

= $300 + $450

= $750

The margin account balance at the end of Day 3 would be $750.

Question: Explain the components making up this amount, $750.
Answer #1

Initial Margin = $ 450

Trading profit after first day = $ 200

Trading profit after second day= $250

Total profit after second day = $450

Trading loss after 3 day = $150

Net profit after 3rd loss= 450-150 =$300

Total amount with investor is =Initial Margin+ Net profit= 450+300 =&750